# Sim-to-Real Domain Adaptation for Deformation Classification

### v1

@@ Abstract
Deformation detection is vital for enabling accurate assessment and prediction of structural changes in materials, ensuring timely and effective interventions to maintain safety and integrity. Automating deformation detection through computer vision is crucial for efficient monitoring, but it faces significant challenges in creating a comprehensive dataset of both deformed and non-deformed objects, which can be difficult to obtain in many scenarios. In this paper, we introduce a novel framework for generating controlled synthetic data that simulates deformed objects. This approach allows for the realistic modeling of object deformations under various conditions. Our framework integrates an intelligent adapter network that facilitates sim-to-real domain adaptation, enhancing classification results without requiring real data from deformed objects. We conduct experiments on domain adaptation and classification tasks and demonstrate that our framework improves sim-to-real classification results compared to simulation baseline.
@@ II Related Works
The detection of deformations has been a critical area of research within the field of computer vision, essential for applications ranging from structural health monitoring [7] to manufacturing and production quality control [8]. Over time, deformation detection approaches have significantly developed to improve the quality and capability of the proposed methods, leading to enhanced precision, reliability, and speed in identifying structural deformities. Existing approaches can be divided into two main streams: Classical approaches, characterized by their reliance on traditional image processing techniques and heuristic algorithms, and Deep Learning-based approaches. Classical approaches often involve edge detection, feature extraction, and template matching, leveraging geometric and photometric invariants to identify deformations. In [9], Erkan et al., propose the integration of gravimetric and magnetic fields through deformation analysis for near-surface detection problems. Authors in [10] use colour encoding to measure transient 3D deformation. Similarly, in [11], bending deformation is detected through changes in colour ratios, observed by compact colour sensors. While these approaches are foundational and important to addressing the deformation detection problem, they often come with certain drawbacks such as their dependency on manual feature selection and engineering, and their sensitivity to noise. Deep learning-based approaches, on the other hand, leverage the power of neural networks to automatically learn and extract features relevant to deformation. In [12], Zhao et al., present an approach for the deformation prediction method using a deep learning model during the numerical control machining process. In [13], introduce a segmentation-based methodology for automating the detection of surface anomalies, including deformations. Their strategy leveraged deep learning for segmentation and a decision network, enabling precise identification and analysis of irregularities on surfaces.
@@ III-A Data Generation
Blender is an open-source computer graphics, rendering, and simulation environment with a variety of tools that make it ideal for the creation of synthetic visual data of deformed and non-deformed objects. BlendTorch [30] was used to help facilitate the data generation process by integrating the Python scripting environment within Blender into an integrated development environment (IDE). The object selected to be deformed was a pop can as the metallic surface features provide an interesting classification challenge and ease of physical dataset collection. For the synthetic dataset, a computer-aided design (CAD) file was imported and using expert information deformations were applied. A lattice deformation linked to the pop can was used to apply larger deformations to the structure while a hard Stucci displace modifier was used on the surface of the can. These various deformations were all assigned shape keys enabling a Python script to create various procedural interpolations of these deformations. In both the deformed and non-deformed classes, shape keys were used to open the tab and seal of the cans. The material shading was set to mimic the metallic look of real-world pop cans and the outer label was wrapped around the object using a UV map. The scene in Blender was lit using randomized High Dynamic Range Images (HDRIs) given a random rotation to provide varietal realistic lighting. Camera rays that did not hit the object were set to black pixels giving the generated image a black background. Cameras were placed stochastically according to Table I and 6000 images were generated; 3000 deformed, 3000 non-deformed. The images were rendered as 512x512 RGB.
@@ III-C Content-Aware Sim-to-real Network (CASNet)
The encoder $E$ consists of a convolutional layer increasing the channels to 32, followed by 3 residual block layers. Finally, a last convolutional layer is used and expands the channels from 32 to 64. The task of the encoder is to take an input image and capture its features.
The separator $S$ consists of three convolutional layers with 64 channels. The purpose of the separator is to split the features from the encoder into content and style feature layers.
The generator $G$ consists of a transpose convolutional layer from 64 to 32, 3 residual block layers and a last transpose convolution layer from 32 to 3. A $\tanh$ activation function is used on the output. The purpose of the generator is to take content and style feature layers and generate an image.
All layers use ReLU as the activation function and spectral normalization [33].
The discriminator takes in images ($I_{X},I_{Y\rightarrow X}$ for $D_{X}$, $I_{Y},I{X\rightarrow Y}$ for $D_{Y}$) and analyzes if the given image is generated or real. The architecture for the patchGAN discriminator was convolution layers from 3 to 64, to 128, to 256, to 512 and finally down to 1 channel. The architecture uses LeakyReLU as an activation function with a negative slope of 0.2.
The perceptual network $P$ is a pre-trained version of VGG-19 on ImageNet1k. The role of the perceptual network is to analyze the images for content and style and help train the separator network. The perceptual network inputs the reconstructed $I^{\prime}_{X}$, $I^{\prime}_{Y}$ and converted images $I_{X\rightarrow Y}$, $I_{Y\rightarrow X}$ and creates feature maps from several middle layers. The network finds perceptual features to create constraints on the content and style of the images to train the separator network in an unsupervised manner.
@@
$\mathcal{H}_{row}=\sigma_{row}(\mathcal{C}_{X}\cdot\mathcal{C}_{Y}^{\top})=\left[\begin{array}[]{ccc}\mathcal{C}_{11}&\cdots&\mathcal{C}_{1b}\\ \vdots&\ddots&\vdots\\ \mathcal{C}_{b1}&\cdots&\mathcal{C}_{bb}\\ \end{array}\right],\quad\mathcal{C}_{X},\mathcal{C}_{Y}\in\mathbb{R}^{B\times N}$ (1)
@@
$\hat{\mathcal{S}_{Y}}=\mathcal{H}_{row}\mathcal{S}_{Y},\quad\text{where}\quad\mathcal{S}_{Y}\in\mathbb{R}^{B\times N}$ (2)
@@ V Results and Discussion
Comparing the generated images from CycleGAN and CASNet shows that the best performance is from CASNet. CycleGAN produced images that copied the style of the background quite well but had a very negative effect on the content of the soda can. The cans in the images produced very loosely resemble their original look and the reconstruction doesn’t match the original image. CASNet produced images where the content is clear and matches the content of the source images. The backgrounds in the generated images are a decent attempt at recreating the target background but have room for improvement. A possible avenue for improving the background style transfer for CASNet could be replacing the Gram style loss with Sliced Wasserstein Discrepancy (SWD) [36]. Experimentation was done with this style loss but it created issues with the content transferred. Additional work is required to develop a method using SWD as style loss.
@@ VI Conclusion
Geometrical quality assurance is an important aspect of industrial processes. The method shown in this work can be used to bring modern machine learning techniques to quality assurance processes that can be expensive and difficult. This is particularly useful in cases where the object of quality control is difficult to create a dataset, e.g., the object is fragile, difficult to handle, expensive, and visual methods are required. The novelty of this work is the use of CASNet for sim-to-real domain adaptation for the use of deformation classification. In this work, CASNet was shown to be able to effectively adapt synthetic datasets from Blender to real-world distributions allowing for better, more realistic data and improved classification metrics. These adapted datasets have a dramatic improvement to deformation classification metrics and reduce the sim-to-real gap present in synthetic datasets. A strength of this method is that minimal unlabelled real-world data is required which reduces time and cost. Our ongoing research aims to improve the style transfer of CASNet using sliced wasserstein discrepancy.

### v2

@@ Abstract
Deformation detection is vital for enabling accurate assessment and prediction of structural changes in materials, ensuring timely and effective interventions to maintain safety and integrity. Automating deformation detection through computer vision is crucial for efficient monitoring, but it faces significant challenges in creating a comprehensive dataset of both deformed and non-deformed objects, which can be difficult to obtain in many scenarios. In this paper, we introduce a novel framework for generating controlled synthetic data that simulates deformed objects. This approach allows for the realistic modeling of object deformations under various conditions. Our framework integrates an intelligent adapter network that facilitates sim-to-real domain adaptation, enhancing classification results requiring limited real data from deformed objects. We conduct experiments on domain adaptation and classification tasks and demonstrate that our framework improves sim-to-real classification results compared to the simulation baseline. Our code is available here.
@@ II Related Works
The detection of deformations has been a critical area of research within the field of computer vision, essential for applications ranging from structural health monitoring [7] to manufacturing and production quality control [8]. Over time, deformation detection approaches have significantly developed to improve the quality and capability of the proposed methods, leading to enhanced precision, reliability, and speed in identifying structural deformities. Existing approaches can be divided into two main streams: Classical approaches, characterized by their reliance on traditional image processing techniques and heuristic algorithms, and Deep Learning-based approaches. Classical approaches often involve edge detection, feature extraction, and template matching, leveraging geometric and photometric invariants to identify deformations. Erkan et al. propose the integration of gravimetric and magnetic fields through deformation analysis for near-surface detection problems [9]. In [10], Mares et al. use colour encoding to measure transient 3D deformation. Similarly, in [11], bending deformation is detected through changes in colour ratios, observed by compact colour sensors. While these approaches are foundational and important to addressing the deformation detection problem, they often come with certain drawbacks such as their dependency on manual feature selection and engineering, and their sensitivity to noise. Deep learning-based approaches, on the other hand, leverage the power of neural networks to automatically learn and extract features relevant to deformation. In [12], Zhao et al., present an approach for the deformation prediction method using a deep learning model during the numerical control machining process. In [13], introduce a segmentation-based methodology for automating the detection of surface anomalies, including deformations. Their strategy leveraged deep learning for segmentation and a decision network, enabling precise identification and analysis of irregularities on surfaces.
@@ III-A Data Generation
Blender is an open-source computer graphics, rendering, and simulation environment with a variety of tools that make it ideal for the creation of synthetic visual data of deformed and non-deformed objects. BlendTorch [30] was used to help facilitate the data generation process by integrating the Python scripting environment within Blender into an integrated development environment. The object selected to be deformed was a pop can as the metallic surface features provide an interesting classification challenge and ease of physical dataset collection. For the synthetic dataset, a computer-aided design (CAD) file was imported and using expert information deformations were applied. A lattice deformation linked to the pop can was used to apply larger deformations to the structure while a hard Stucci displace modifier was used on the surface of the can. These various deformations were all assigned shape keys enabling a Python script to create various procedural interpolations of these deformations. In both the deformed and non-deformed classes, shape keys were used to open the tab and seal of the cans. The material shading was set to mimic the metallic look of real-world pop cans and the outer label was wrapped around the object using a UV map. The scene in Blender was lit using randomized High Dynamic Range Images (HDRIs) given a random rotation to provide varietal realistic lighting. Camera rays that did not hit the object were set to black pixels giving the generated image a black background. Cameras were placed stochastically according to Table I and 6000 images were generated; 3000 deformed, 3000 non-deformed. The images were rendered as 512x512 RGB.
@@ III-C Content-Aware Sim-to-real Network (CASNet)
The encoder $E$ consists of a convolutional layer increasing the channels to 32, followed by 3 residual block layers. Finally, a last convolutional layer is used and expands the channels from 32 to 64. The task of the encoder is to take an input image and capture its features. The separator $S$ consists of three convolutional layers with 64 channels. The purpose of the separator is to split the features from the encoder into content and style feature layers. The generator $G$ consists of a transpose convolutional layer from 64 to 32, 3 residual block layers and a last transpose convolution layer from 32 to 3. A $\tanh$ activation function is used on the output. The purpose of the generator is to take content and style feature layers and generate an image. All layers use ReLU as the activation function and spectral normalization [33]. The discriminator takes in images ($I_{X},I_{Y\rightarrow X}$ for $D_{X}$, $I_{Y},I{X\rightarrow Y}$ for $D_{Y}$) and analyzes if the given image is generated or real. The architecture for the patchGAN discriminator was convolution layers from 3 to 64, to 128, to 256, to 512 and finally down to 1 channel. The architecture uses LeakyReLU as an activation function with a negative slope of 0.2. The perceptual network $P$ is a pre-trained version of VGG-19 on ImageNet1k. The role of the perceptual network is to analyze the images for content and style and help train the separator network. The perceptual network inputs the reconstructed $I^{\prime}_{X}$, $I^{\prime}_{Y}$ and converted images $I_{X\rightarrow Y}$, $I_{Y\rightarrow X}$ and creates feature maps from several middle layers. The network finds perceptual features to create constraints on the content and style of the images to train the separator network in an unsupervised manner.
@@
$\begin{split}\mathcal{H}_{row}=\sigma_{row}(\mathcal{C}_{X}\cdot\mathcal{C}_{Y}^{\top})=\left[\begin{array}[]{ccc}\mathcal{C}_{11}&\cdots&\mathcal{C}_{1b}\\ \vdots&\ddots&\vdots\\ \mathcal{C}_{b1}&\cdots&\mathcal{C}_{bb}\\ \end{array}\right],\\ \text{where}\ \mathcal{C}_{X},\mathcal{C}_{Y}\in\mathbb{R}^{B\times N}\end{split}$ (1)
@@
$\hat{\mathcal{S}_{Y}}=\mathcal{H}_{row}\mathcal{S}_{Y},\quad\text{where}\ \mathcal{S}_{Y}\in\mathbb{R}^{B\times N}$ (2)
@@ V Results and Discussion
Comparing the generated images from CycleGAN and CASNet shows that the best performance is from CASNet. CycleGAN produced images that copied the style of the background quite well but had a very negative effect on the content of the soda can. The cans in the images produced very loosely resemble their original look and the reconstruction doesn’t match the original image. CASNet produced images where the content is clear and matches the content of the source images. The backgrounds in the generated images are a decent attempt at recreating the target background but have room for improvement. A possible avenue for improving the background style transfer for CASNet could be replacing the Gram style loss with Sliced Wasserstein Discrepancy (SWD) [36]. Experimentation was done with SWD style loss but results had issues where the content sometimes failed to transfer. Overall the style transfer was improved with this loss but additional work is required to ensure proper content transfer.
@@ VI Conclusion
Geometrical quality assurance is an important aspect of industrial processes. The method shown in this work can be used to bring modern machine learning techniques to quality assurance processes that can be expensive and difficult. This is particularly useful in cases where the object of quality control is difficult to create a dataset, e.g., the object is fragile, difficult to handle, expensive, and visual methods are required. The novelty of this work is the use of CASNet for sim-to-real domain adaptation for the use of deformation classification. In this work, CASNet was shown to be able to effectively adapt synthetic datasets from Blender to real-world distributions allowing for better, more realistic data and improved classification metrics. These adapted datasets have a dramatic improvement to deformation classification metrics and reduce the sim-to-real gap present in synthetic datasets. A strength of this method is that minimal unlabelled real-world data is required which reduces time and cost. Our ongoing research aims to improve the style transfer of CASNet using sliced Wasserstein discrepancy. In future work, we plan to incorporate depth information into the datasets as it represents a promising avenue for enhancing model performance. Depth data provides a critical spatial understanding that can significantly improve the accuracy and robustness of the model, especially for tasks requiring precise object localization and scene comprehension. Additionally, exploring different network architectures such as vision transformers could yield better performance.